### v6

@@ 3.4 Innovating Spatial Attention and Standard Convolutional Operation
Figure 7: The detailed structure of RFAConv, which dynamically determines the importance of each feature in the receptive-field and solves the problem of parameters sharing.
As one of the popular frameworks in the field of deep learning, Pytorch provides the Unfold method to extract receptive-field spatial features. The detailed structure is shown in Fig. 6, which extracts 3$\times$3 receptive-field spatial features. Let input X $\in$ $R^{C\times H\times W}$, after the Unfold method, its dimension becomes 9C×$\times$×H×$\times$×W. Where C, H, and W represent the number of channels, height and width of input. Although Unfold is able to extract receptive-field spatial features by parameterless way, it is slow. Therefore, in RFAConv, we utilize a fast method to extract receptive-field spatial features, i.e., Group Conv. As mentioned in the previous section, each 3$\times$3 size window in the receptive-field spatial feature represents a receptive-field slider when the 3$\times$3 convolution kernel is used to extract features. However, after using fast Group Conv to extract the receptive-field features, original features are mapped into new features. This method is fast and more efficient than the original Unfold method. As shown in Table 1, experiments based on the YOLOv5n and VisDrone dataset [40] demonstrate it. It can be seen that RFAConv based on GroupConv obtains good performance while training 300 epochs requires less training time than the Unfold method. Morefore, we need to explain that the Unfold method is parameterless, while in the Table 1, it can be seen that the number of parameters required for the GroupConv-based method is the same as for the Unfold method. Because we use a lightweight approach to interacting information in receptive-field.
@@
Recent research has shown that interacting information can enhance network performance, as demonstrated in [41, 42, 43]. Similarly, for RFAConv, interacting receptive-field feature information to learn the attention map can enhance network performance. However, interacting with each receptive-field feature can result in additional computational overhead,so to minimize computational overhead and number of parameters, the AvgPool is utilized to aggregate the global information of each receptive-field feature. Then, a 1×1 group convolutional operation is used to interact information. Finally, we use softmax to emphasize the significance of each feature within the receptive-field feature. In general, the calculation of RFA can be expressed as:
@@
Here, $g^{i\times i}$ represents a grouping convolution of size $i\times i$, k represents the size of the convolution kernel, Norm stands for normalization, X represents the input feature maps, F is obtained by multiplying the attention map $A_{rf}$ with the transformed receptive-field spatial feature $F_{rf}$. Unlike CBAM and CA, RFA is capable of generating attention maps for each receptive-field feature. The performance of convolutional neural networks is limited by the standard convolutional operations due to the fact that the convolution operation relies on shared parameters and is not sensitive to differences in information brought about by positional variations. However, RFAConv can completely address this issue by emphasizing the significance of different features within the receptive field slider and prioritizing the receptive-field spatial feature.
The feature map obtained through RFA is the receptive-field spatial feature, which does not overlap after ”Adjust Shape”. Therefore, the learned attention map aggregate the feature information of each receptive field slider. In other words, the attention map is no longer shared within each receptive-field slider. This completely compensates for the shortcomings of the existing CA and CBAM attention mechanisms. RFA provides significant benefits to the standard convolution kernel. However, after adjusting shape, the features are k times in height and width, requiring a k $\times$ k convolution operation with a stride = k to extract feature information. The convolution operation RFAConv designed by RFA brings good gains for convolution, and it innovates the standard convolution.
@@
Furthermore, We assert that existing spatial attention mechanisms, such as CBAM [17] and CA [18], should prioritize receptive field spatial features to improve network performance. As is well-known, the network model based on the self-attention mechanism [44, 45, 46] has achieved great success, because it solves the problem of convolution parameter sharing and models long-distance information. However, the self-attention mechanism also introduces significant computational overhead and complexity to the model. We argue that directing the attention of existing spatial attention mechanisms to the receptive-field spatial feature can solve the problems of parameter sharing and modeling of long-range information in a way similar to self-attention. This approach requires significantly fewer parameters and computational resources than self-attention. The answers are as follows:
Therefore, we design new CBAM and CA models called RFCBAM and RFCA, which focus on the receptive-field spatial feature. Similar to RFA, the final convolution operation of $k\times k$ with stride = k is used to extract the feature information. The specific structure of these two new convolution methods, as shown in Fig. 8, we call these two new convolution operations RFCBAMConv and RFCAConv. Comparing the original CBAM, We use SE attention to replace CAM in RFCBAM. Because this can reduce computational overhead. Moreover in RFCBAM, channel and spatial attention are not performed in separate steps. Instead, they are weighted simultaneously, allowing the attention map obtained for each channel to be different.
@@ 1 $\times$ 1        AvgPool 1000-d
4 Experiments and Discussions
To verify the effectiveness of our method, we perform classification, object detection and semantic segmentation experiments. The equipment for all experiments are based on RTX3090. In the classification experiments, we use four RTX3090 to train the model in parallel.
4.1 Classification experiments on ImageNet-1k
We perform experiments to validate our method on the ImageNet-1k, which contains a number of 1281167 training sets and 50000 validation sets. Similar to RFAConv, we construct CBAMConv and CAConv by combining the CBAM and CA, respectively, with an additional 3 $\times$ 3 convolution layer at the end of the CBAM and CA modules. We also compare the CAMConv, which is constructed using the channel attention mechanism CAM [17]. we evaluate in ResNet18, ResNet34. Specifically, RFAConv, CBAMConv, CAConv, and CAMConv are used to replace (r) the first convolutional layer of BasicBlock in ResNet18 and ResNet34, respectively. In general, the new convolution is structured as shown in Table 2.
@@ 4.1 Classification experiments on ImageNet-1k
Table 3: Classification results on ImageNet-1K using the Resnet18 and Resnet34. The different convolutional operation constructed by the attention mechanism is compared.
The NewConv in Table 2 represents the convolution mode constructed by the attention mechanism. For CAConv, CBAMConv, and CAMConv, it can be considered that CA, CBAM, and CAM are added before the first layer of convolution in BasicBlock. In the image classification experiments, we train 100 epochs for each model with batch-size of 128. The learning rate start from 0.1 and decrease every 30 epochs, 0.1 times each time. In experiments, we follow most of the previous work and report the accuracy for TOP1 and TOP5, respectively. Table 3 shows the results produced by different networks on the ImageNet-1K validation set. It is clear that replacing the 3$\times$3 convolutional operation with RFAConv significantly improves the recognition results. Compared to the baseline models RestNet18 and ResNet34 , the network constructed by RFAConv achieves the best recognition results at the cost of only a small increase in parameters and computational overhead. Such as ResNet18 constructed based on RFAConv only adds 0.16 M parameters and 0.09 G computational overhead over the original model, and increases the accuracy by 1.64% and 1.24% on TOP1 and TOP5, respectively.
Moreover, as we mentioned earlier, spatial attention can be enhanced again by focusing on receptive-field spatial features. Therefore, we design RFCBAMConv and RFCAConv, which are improvements of CBAM and CA. In order to verify their advantages, we conduct experiments based on ResNet18 and report the relevant data in Table 4. It is obvious that RFCBAMConv and RFCAConv achieve better recognition accuracy, compared with CBAMConv and CAConv in Table 3. Most importantly, they also significantly improve performance at the cost of only a small increase in parameters and computational overhead. This is a strong demonstration that spatial attention can be improved by placing attention into the receptive-field spatial features. It fully demonstrates that spatial attention can be improved by placing attention into the receptive-field spatial features.
@@
+ RFCAConv(r) 1.92 11.89 72.01 90.64
Table 4: RFCBAMConv and RFCAConv improve the performance of CBAMConv and CAConv. The table shows that the classification accuracy is significantly improved on ImageNet-1k.
@@
Moreover, to provide a more intuitive analysis, as with the most work, we use the Grad-CAM [47] algorithm for visualization. Grad-CAM highlights the regions of interest of different networks for a particular class of objects. To some extent, it is possible to see how the networks utilize the features. We randomly selected some images in the validation set of ImageNet-1K and visualized the results of the networks constructed with different attention convolutions based on ResNet18 separately. As shown in the Fig. 9, compared with other attention convolutions, our RFAConv can help the network to better recognize and highlight the key regions of objects.
@@
We also use Grad-CAM to visualize ResNet18 constructed by CBAMConv, RFCBAMConv, CAConv, and RFCAConv. RFCBAM is obtained by improving the attention of CBAM by putting it into the recepitve-field spatial feature. Similarly, RFCA is obtained by CA after the same method. As shown in Fig. 10, the improved RFCA and RFCBAM can help the network to better recognize and highlight the key regions of objects after combining with the convolution operation.
@@
Figure 10: We put the attention of CBAM and CA into the receptive-field spatial features and improve them to obtain RFCBAM and RFCA. Then RFCBAMConv and RFCAConv are constructed by the same method as RFA. As in Fig. 10, we visualize the different networks separately. Obviously, compared to CBAMConv and CAConv, the improved obtained RFCBAMConv and CAConv can help the network to better recognize and highlight the key regions of objects.
@@ 4.2 Object detection experiments on COCO2017
We conduct object detection experiments on COCO2017 to re-evaluate our methods. COCO2017 contains 118287 training sets and 5000 verification sets. We select YOLOv5n, YOLOv7-tiny, and YOLOv8n models to perform a series of experiments. All parameters except for epoch and batch-size are set to the default values. We train each model for 300 epochs with a batch-size of 32. To be similar in classification, we replace some convolution operation in the baseline model with novel convolution operations constructed using attention mechanisms. Specifically, we replace all 3$\times$3 convolution operations in the yaml files for YOLOv5 and YOLOv8 using attention convolution. And for YOLOv7, we replace the first 3$\times$3 convolution operation in all ELAN [34] in the backbone.
@@
+ CAConv(r) 4.5 1.8 46.2 29.2 28.1 14.3 32 36.6 4.8
+ RFAConv(r) 4.7 1.9 47.3 30.6 29 14.8 33.4 37.4 5.3
YOLOv7-tiny 13.7 6.2 53.8 38.3 35.9 19.9 39.4 48.8 6.8
+ RFAConv(r) 14.1 6.3 55.1 40.1 37.1 20.9 41.1 50 8.4
YOLOv8n 8.7 3.1 51.9 39.7 36.4 18.4 40.1 52 4.2
+ CAMConv(r) 8.8 3.1 51.6 39 36.2 18 39.9 51.2 4.5
@@
Table 5: Object detection $AP_{50}$, $AP_{75}$, $AP$, $AP_{S}$, $AP_{M}$, and $AP_{L}$ on the COCO2017 validation sets. We adopt the YOLOv5n, YOLOv7-tiny, and YOLOv8n detection framework and replace the original convolution with the novel convolutioanl operation constructed by attention mechanism.
Following the previous work, we report $AP_{50}$, $AP_{75}$, $AP$, $AP_{S}$, $AP_{M}$ and $AP_{L}$ separately. Moreover, in order to better display the performance of different networks, we chose the training process of YOLOv5n for visualization. We visualize how the AP50 changes with the number of iterations. The experimental results are shown in Table 5 and Fig. 11. When RFAConv is used to replace some convolution, the network achieved significantly improved detection results with only a small increase in the number of parameters and computational overhead. Compared to other attention, RFA still brings considerable benefits to the detection network. In some experiments, we once again verify the effectiveness of RFCA, which exhibits better performance of convolutional operation compared to the original CA. Time represents the total time spent processing an image during validation. It can be clearly seen that the model constructed with the novel convolutioanl operation has an increase in time for processing a image. Therefore, if real-time is pursued, the number of replacement convolutions should not be too many.
Figure 11: The $AP_{50}$ change during training for the different YOLOv5n constructed by attention convolution.

### v7

@@ 3.3 Receptive-Filed Attention Convolution
Figure 4: The detailed structure of RFAConv, which dynamically determines the importance of each feature in the receptive-field and solves the problem of parameters sharing. In the receptive-field spatial feature, the nine features adjacent to each channel represent the features in receptive-field slider at the original input corresponding position. Therefore, it needs to be adjusted to the spatial dimension by "Adjust shape" to facilitate the subsequent convolution operation.
@@
Figure 5: In the figure, it shows in detail an example of extracting 3 $\times$ 3 receptive-field spatial features by the Unfold method. It is worth noting that the unfold method performs the same zero-padding process on the feature map as the convolution, so the final output has the same width and height as the input. In the figure, we ignore the zero-padding for the convenience of showing the process of the unfold method.
Methods mAP50($\%$) mAP($\%$) FLOPS(G) Param(M) Training Time (Hours)
@@
Receptive-Field Attention Process: Recent research has shown that interacting information can enhance network performance, as demonstrated in [22]. Similarly, for RFAConv, interacting receptive-field feature information to learn the attention map can enhance network performance. However, interacting with each receptive-field feature can result in additional computational overhead, so to minimize computational overhead and number of parameters, the AvgPool is utilized to aggregate the global information of each receptive-field feature. Then, a 1 $\times$ 1 group convolutional operation is used to interact information. Finally, we use softmax to emphasize the significance of each feature within the receptive-field feature. The range of features targeted by softmax are the features of the corresponding receptive-field slider on each channel, which is analogous to grouping features by the number of channels. In general, the calculation of RFA can be expressed as:
@@
Here, $g^{i\times i}$ represents a grouping convolution of size $i\times i$, k represents the size of the convolution kernel, Norm stands for normalization, $X$ represents the input feature maps, $F$ is obtained by multiplying the attention map $A_{rf}$ with the transformed receptive-field spatial feature $F_{rf}$. Unlike traditional spatial attention, such as CBAM and CA, RFA assigns different weights to each feature in the receptive-field slider by softmax operation, and the weights obtained by learning are not shared in each convolutional slider after extension, as shown in Figure 6. Attentional weights learned by traditional spatial attention mechanisms are shared across each receptive-field slider. In contrast, the attentional weights learned by the RFA do not overlap, so each receptive-field slider has independent attentional weights that do not share weights from other sliders. The performance of convolutional neural networks is limited by the standard convolutional operations due to the fact that the convolution operation relies on shared parameters and is not sensitive to differences in information brought about by positional variations. However, RFAConv can completely address this issue by emphasizing the significance of different features within the receptive-field slider and prioritizing the receptive-field spatial feature. The feature map obtained through RFA is the receptive-field spatial feature, which does not overlap after "Adjust Shape" (It transforms the receptive-field slider stacked on the channel dimension into the spatial dimension to complete the subsequent convolution operation).
Figure 6: (a) and (b) denote the attentional weights obtained by traditional spatial attention and receptive-field attention, respectively. It is evident that traditional spatial attention does not unfold the receptive-slider features, which leads to problems with sharing the obtained attentional weights.
After the two-step process described above, the features are k times in height and width, requiring a k $\times$ k convolution operation with a stride = k to extract feature information. The process of extracting features in the receptive-field spatial feature by a k $\times$ k convolution operation with a stride = k just corresponds to the process of convolutional operation in original spatial features. RFA focuses on the receptive-field spatial feature, so the weight of the attention gained is not shared among each receptive-field slider. Therefore, when we consider the last convolution operation as a whole with the previous attention weighting operation, the problem of convolution parameter sharing is perfectly solved. Be careful not to look at the subsequent convolution operation separately, but to see the attention weighting operation and the convolution operation as a whole. The convolution operation RFAConv designed by RFA brings good gains for convolution, and it innovates the standard convolution.
Furthermore, We assert that existing spatial attention mechanisms, such as CBAM [6] and CA [7], should prioritize receptive-field spatial features to improve network performance. As is well-known, the network model based on the self-attention mechanism [23] has achieved great success, because it solves the problem of convolution parameter sharing and models long-distance information. However, the self-attention mechanism also introduces significant computational overhead and complexity to the model. We argue that directing the attention of existing spatial attention mechanisms to the receptive-field spatial feature can solve the problems of parameter sharing and modeling of long-range information in a way similar to self-attention. This approach requires significantly fewer parameters and computational resources than self-attention. The answers are as follows:
Therefore, we design new CBAM and CA models called RFCBAM and RFCA, which focus on the receptive-field spatial feature. Similar to RFA, the final convolution operation of $k\times k$ with stride = k is used to extract the feature information. The specific structure of these two new convolution methods, as shown in Figure 7, we call these two new convolution operations RFCBAMConv and RFCAConv. Comparing the original CBAM, We use Squeeze-and-Excitation attention to replace CAM in RFCBAM. Because this can reduce computational overhead. Moreover in RFCBAM, channel and spatial attention are not performed in separate steps. Instead, they are weighted simultaneously, allowing the attention map obtained for each channel to be different.
@@
4 Experiments and Discussions
To verify the effectiveness of our method, we perform classification, object detection and semantic segmentation experiments. The equipment for all experiments are based on RTX3090. In the classification experiments on ImageNet-1k, we use four RTX3090 to train the model in parallel. In the experiments, Cuda version and Pytroch version are 11.1 and 1.8.1, respectively.
4.1 Classification experiments on ImageNet-1k
Currently, for classification tasks, ImageNet-1k is a large-scale dataset that contains 1.28 million training sets, and 50,000 validation sets. Therefore, We perform experiments to validate our method on the ImageNet-1k. Similar to RFAConv, we construct CBAMConv and CAConv by combining the CBAM and CA, respectively, with an additional 3 $\times$ 3 convolution layer at the end of the CBAM and CA modules. We also compare the CAMConv, which is constructed using the channel attention mechanism CAM [6]. we evaluate in ResNet18, ResNet34. Specifically, RFAConv, CBAMConv, CAConv, and CAMConv are used to replace (r) the first convolutional layer of BasicBlock in ResNet18 and ResNet34, respectively. In general, the new convolution is structured as shown in Table 2.
@@ 1 $\times$ 1 AvgPool 1000-d
The NewConv in Table 2 represents the convolution mode constructed by the attention mechanism. For CAConv, CBAMConv, and CAMConv, it can be considered that CA, CBAM, and CAM are added before the first layer of convolution in BasicBlock. In the image classification experiments, we train 100 epochs for each model with batch-size of 128. The learning rate start from 0.1 and decrease every 30 epochs, 0.1 times each time. In experiments, we follow most of the previous work and report the accuracy for TOP1 and TOP5, respectively. Table 3 shows the results produced by different networks on the ImageNet-1K validation set. It is clear that replacing the 3 $\times$ 3 convolutional operation with RFAConv significantly improves the recognition results. Compared to the baseline models RestNet18 and ResNet34 , the network constructed by RFAConv achieves the best recognition results at the cost of only a small increase in parameters and computational overhead. Such as ResNet18 constructed based on RFAConv only adds 0.16 M parameters and 0.09 G computational overhead over the original model, and increases the accuracy by 1.64% and 1.24% on TOP1 and TOP5, respectively. Whether it is RFAConv, CBAMConv, CAConv, or other convolutional operations constructed for attention mechanisms, it can be seen as adding attention mechanisms in front of convolution. So to compare the work of other attention mechanisms added to ResNet, we also compare ECA [16], QEA [24], QCA [25], EMAS [26], TA [18] and GAM [27] in Table 3. Compared to these works, RFAConv can still get good performance.
Table 3: Classification results on ImageNet-1K using the ResNet18 and ResNet34. The different convolutional operation constructed by the attention mechanism is compared. Moreover, some related works are compared.
@@
Moreover, as we mentioned earlier, spatial attention can be enhanced again by focusing on receptive-field spatial features. Therefore, we design RFCBAMConv and RFCAConv, which are improvements of CBAM and CA. In order to verify their advantages, we conduct experiments based on ResNet18 and report the relevant data in Table 4. It is obvious that RFCBAMConv and RFCAConv achieve better recognition accuracy, compared with CBAMConv and CAConv in Table 3. Most importantly, they also significantly improve performance at the cost of only a small increase in parameters and computational overhead. This is a strong demonstration that spatial attention can be improved by placing attention into the receptive-field spatial features. It fully demonstrates that spatial attention can be improved by placing attention into the receptive-field spatial features. In the Table 4, we have also compared other excellent works. Compared to CBAM, works such as AFF-CAM [28], MCA [19], LMA [29] and DAS [30] have been able to obtain satisfactory results for the network. But the performance obtained from RFCBAM obtained by focusing on receptive-field spatial features to improve CBAM and combining convolutional operations has obtained better performance.
Table 4: RFCBAMConv and RFCAConv improve the performance of CBAMConv and CAConv, and significantly improve the accuracy obtained by the network compared to other excellent works. The table shows that the classification accuracy is significantly improved on ImageNet-1k.
@@
Moreover, to provide a more intuitive analysis, as with the most work, we use the Grad-CAM [31] algorithm for visualization. Grad-CAM highlights the regions of interest of different networks for a particular class of objects. To some extent, it is possible to see how the networks utilize the features. We randomly selected some images in the validation set of ImageNet-1K and visualized the results of the networks constructed with different attention convolutions based on ResNet18 separately. As shown in the Figure 8, compared with other attention convolutions, our RFAConv can help the network to better recognize and highlight the key regions of objects.
@@
Figure 9: We put the attention of CBAM and CA into the receptive-field spatial features and improve them to obtain RFCBAM and RFCA. Then RFCBAMConv and RFCAConv are constructed by the same method as RFA. Obviously, compared to CBAMConv and CAConv, the improved obtained RFCBAMConv and CAConv can help the network to better recognize and highlight the key regions of objects.
Figure 10: It shows the accuracy change curves of ResNet18 for TOP1 and TOP5 on the validation set during the training process.
We also use Grad-CAM to visualize ResNet18 constructed by CBAMConv, RFCBAMConv, CAConv, and RFCAConv. RFCBAM is obtained by improving the attention of CBAM by putting it into the receptive-field spatial feature. Similarly, RFCA is obtained by CA after the same method. As shown in Figure 9, the improved RFCA and RFCBAM can help the network to better recognize and highlight the key regions of objects after combining with the convolution operation. Moreover, to further observe the accuracy changes during the training process of networks constructed by different attention convolutions, we visualized the TOP1 and TOP5 data obtained by ResNet18 on the validation set. As shown in Figure 10, it is evident that compared to other methods, convolutional operations focusing on receptive-field spatial features can consistently maintain a higher level of accuracy in the network.
4.2 Classification experiments on ImageNet-200 and Places365
In order to verify that the proposed method can improve the performance of different networks, we conduct corresponding experiments based on the ImageNet-200 dataset. Compared with ImageNet-1k, ImageNet-200 contains 100000 training sets and 10000 validation sets, which can again verify whether the proposed method can gain advantages even with small amount of data. In experiments, we select six networks of different sizes for the corresponding classification experiments, such as FasterNet [32], InceptionNeXt [33], MobileNet [34], ShuffleNet [12]. For these networks, we utilize the proposed convolution operation to replace the large-size standard convolution operation in their networks for comparison experiments. In the experiment, all the parameter settings are similar to the previous section. Table 5 shows the corresponding experimental results. Compared with the baseline model, after replacing the standard convolution operation with the proposed convolution operation, the classification performance is better than the baseline model, despite the increase in the number of small parameters and computational overhead. Therefore, whether it is the large-scale dataset ImageNet-1K or the smaller dataset ImageNet-200, our proposed methods are able to obtain good performance. The reason is that different images exhibit varying information disparities at different positions, yet the parameter-sharing mechanism of standard convolution fails to capture these distinct information patterns. In contrast, the proposed convolution operation can employ distinct convolution kernels at different spatial positions, thereby adaptively capturing these location-specific information disparities.
Table 5: Classification results on ImageNet-200 using different baseline model. Different networks constructed by the RFCAConv are compared.
Models FLOPS(G) Param(M) Top1(%) Top5(%)
FasterNet 157.53 13.96 41.28 67.88
FasterNet + RFCAConv(r) 161.58 14.12 41.84 68.17
InceptionNeXt 685.42 47.52 20.26 45.01
InceptionNeXt + RFCAConv 690.98 47.73 28.03 55.00
MobileNet 2.79 0.61 36.26 63.56
MobileNet + RFCAConv(r) 5.48 0.72 40.28 67.78
ShuffleNet 12.22 1.11 42.23 68.66
ShuffleNet + RFCAConv(r) 18.11 1.38 43.68 69.64
Moreover, to further compare other pure attention-based architectures and hybrid architectures, we construct a corresponding network based on InceptionNeXt by replacing some large-sized convolutions with RFCAConv, and conduct corresponding experiments on the large-scale scene recognition dataset Places365. Places365 contains a total of 365 categories, with over 1.8 million images in total. Table 6 presents the corresponding experimental results, which are basically consistent with the results in [35]. The results obtained by InceptionNeXt-A [33] are similar to those achieved by the pure attention mechanism architecture. However, in the Places365 dataset, the classification accuracy of InceptionNeXt-A is lower compared to other networks. This is understandable, as Places365 is a scene classification dataset that requires more contextual information. Therefore, compared to other attention mechanism networks, InceptionNeXt-A achieves lower accuracy. However, after incorporating RFCAConv into InceptionNeXt-A, our network achieves better results than PVTv2-B0 [36], DeiT-T [37], and T2T-ViT-7 [38] under nearly identical parameters and computational overheads.
Table 6: Results of classification using different network architectures on the Places365 dataset.
Models FLOPS(G) Param(M) Top1(%) Top5(%)
PVTv2-B0 0.53 3.5 34.25 64.28
DeiT-T 1.07 5.56 36.33 66.72
T2T-ViT-7 0.98 4.09 37.56 68.75
InceptionNeXt-A 0.5 3.54 35.67 65.78
Ours 0.52 3.6 38.14 69.1
@@ 4.3 Object detection experiments on COCO2017
We conduct object detection experiments on COCO2017 to re-evaluate our methods. COCO2017 contains 118287 training sets and 5000 verification sets. We select YOLOv5n, YOLOv7-tiny, and YOLOv8n models to perform a series of experiments.
All parameters except for epoch and batch-size are set to the default values. We train each model for 300 epochs with a batch-size of 32. To be similar in classification, we replace some convolution operation in the baseline model with novel convolution operations constructed using attention mechanisms, such as CAM, CBAM, and CA. Moreover, to fairly compare the performance of TA [18], HAM [20], MCA [19], the same convolutional operations are added behind these attention mechanisms to constitute TAConv, MCAConv and HAMConv. Similarly, we also use them to replace the convolutional operations at the same locations to conduct related experiments. Specifically, we replace all 3 $\times$ 3 convolution operations in the yaml files for YOLOv5 and YOLOv8 using attention convolution. And for YOLOv7, we replace the first 3 $\times$ 3 convolution operation in all ELAN [14] in the backbone. Note that HAMConv replacing the convolution operation at the same location in YOLOv8n would create a NaN problem in training, so we do not use HAMConv in YOLOv8n.
Following the previous work, we report $AP_{50}$, $_{75}, APAP$, $AP_{S}$, $AP_{M}$ and $AP_{L}$ separately. Moreover, in order to better display the performance of different networks, we chose the training process of detection network for visualization. We visualize how the AP50 and AP changes with the number of iterations. The experimental results are shown in Table 7 and Figure 11. When RFAConv is used to replace some convolution, the network achieved significantly improved detection results with only a small increase in the number of parameters and computational overhead. Compared to other attention, RFA still brings considerable benefits to the detection network. In experiments, we once again verify the effectiveness of RFCBAM and RFCA, which exhibits better performance of convolutional operation compared to the original CBAM and CA. This is precisely thanks to the consideration of receptive-field spatial features to solve the problem of convolutional parameter sharing. Time represents the total time spent processing an image during validation. It can be clearly seen that the model constructed with the novel convolutioanl operation has an increase in time for processing a image. Therefore, if real-time is pursued, the number of replacement convolutions should not be too many.
Table 7: Object detection $AP_{50}$, $_{75}, APAP$, $AP_{S}$, $AP_{M}$, and $AP_{L}$ on the COCO2017 validation sets. We adopt the YOLOv5n, YOLOv7-tiny, and YOLOv8n detection framework and replace the original convolution with the novel convolutioanl operation constructed by attention mechanism.
@@
+ CAConv(r) 4.5 1.8 46.2 29.2 28.1 14.3 32.0 36.6 4.8
+ TAConv(r) 4.8 1.9 46.3 29.0 27.9 13.9 31.5 36.5 5.5
+ HAMConv(r) 4.6 1.9 45.7 28.2 27.4 15.0 31.3 35.6 5.7
+ MCAConv(r) 4.5 1.9 45.3 28.6 27.3 13.8 31.0 36.0 5.5
+ RFAConv(r) 4.7 1.9 47.3 30.6 29.0 14.8 33.4 37.4 5.3
+ RFCBAMConv(r) 4.7 1.9 48.2 30.8 29.4 15.2 33.9 38.0 5.4
+ RFCAConv(r) 4.8 1.9 48.0 30.9 29.5 15.6 33.3 38.2 5.5
YOLOv7-tiny 13.7 6.2 53.8 38.3 35.9 19.9 39.4 48.8 3.4
+ TAConv(r) 13.9 6.2 54.6 39.6 36.7 20.5 41.1 49.1 4.1
+ HAMConv(r) 13.7 6.2 53.8 38.4 35.9 20.4 39.6 48.8 4.2
+ MCAConv(r) 13.7 6.2 53.9 38.9 36.2 20.2 40.1 48.8 3.9
+ RFAConv(r) 14.1 6.3 55.1 40.1 37.1 20.9 41.1 50 4.4
+ RFCBAMConv(r) 14.1 6.3 55.2 40.0 37.1 20.1 41.1 50.4 4.6
+ RFCAConv(r) 14.2 6.3 55.1 40.0 37.1 20.8 41.0 49.5 4.6
YOLOv8n 8.7 3.1 51.9 39.7 36.4 18.4 40.1 52.0 4.2
+ CAMConv(r) 8.8 3.1 51.6 39.0 36.2 18 39.9 51.2 4.5
@@
Figure 11: The $AP_{50}$ and APAP change during training for the different YOLOv5n, YOLOv7-tiny and YOLOv8n constructed by attention convolution.
4.4 Object detection experiments on VOC7+12 and Roboflow-100
Table 8: Object detection mAP50 and mAP on the VOC7+12 validation set.